can you help me solve this: [18(a^5)(b^5)(c^10)] / [15(a^5)(c^15)]
thank you!!

Answer by checkley75(3666)

[18a^5b^5c^10]/[15a^5c^15] reduce the numbers 18/15=6/5, cancel out the a^5 &
reduce c^10/c^15=1/c^5 thus we have
[6b^5]/[5c^5]